# Social Mechanics

### 20 Feb 2024 v1

@@ 3.1 The individual’s motion
$F=m\,\frac{{\rm d}^{2}}{{\rm d}t^{2}}x+\left(\frac{{\rm d}}{{\rm d}x}m\right)\left(\frac{{\rm d}}{{\rm d}t}x\right)^{2}\,.$ (3)
Therefore, a key proposal of this work, in terms of a mathematical formalism, is that individual evolution might be modelled as particles with position-dependent masses.
@@
$0=x^{2}\,\frac{{\rm d}^{2}}{{\rm d}t^{2}}x+2x\left(\frac{{\rm d}}{{\rm d}t}x\right)^{2}\,.$ (5)
@@ 3.3 Stochastic motion
where, in general, $\alpha=\alpha(x,t)$ and $\beta=\beta(x,t)$ are called the diffusion and drift, respectively. This equation is used to explain the time evolution of the probability density function of a particle under the influence of random and drag forces, described by $\alpha$ and $\beta$, respectively. In a social mechanical sense, this model describes a population subject to a compelling trend $\beta$ and some noise $\alpha$ that scatters agglomerations of beliefs. For a simple case, the result is a non-uniform distribution that moves in time changing both its mean, as well as its variance. Let us see how.
@@
$\alpha=\frac{\omega_{\beta}}{2}\sin(\omega_{\alpha}t)\,,\qquad\beta=-\omega_{\beta}\sin(\omega_{\beta}t)\,.$ (14)
@@
Admittedly, this is a rather simplified model of the changes in political stances in the USA, but a very descriptive initial one, nonetheless. This model also shows that the framework does not explain the nature of the interactions. Although, it may shed some light on certain processes and, more specifically, on effective phenomena.

### 27 Feb 2024 v2

@@ 3.1 The motion of a particle
$F=m\,\frac{{\rm d}^{2}}{{\rm d}t^{2}}x+\left(\frac{{\rm d}}{{\rm d}x}m\right)\left(\frac{{\rm d}}{{\rm d}t}x\right)^{2}\,,$ (3)
where we have used de notation ${\rm d}_{t}=\frac{{\rm d}}{{\rm d}t}$ and ${\rm d}_{x}=\frac{{\rm d}}{{\rm d}x}$. We shall use this notation from now on, as well as its partial derivative forms $\partial_{t}$ and $\partial_{x}$. Therefore, a key proposal of this work, in terms of a mathematical formalism, is that individual evolution might be modelled as particles with position-dependent masses.
@@
$0=x^{2}\,{\rm d}_{t}^{2}x+2x\left({\rm d}_{t}x\right)^{2}\,.$ (5)
@@ 3.3 Stochastic motion
where, in general, $\alpha=\alpha(x,t)$ and $\beta=\beta(x,t)$ are called the diffusion and drift, respectively. This equation can be obtained from the overdamped Brownian motion of a particle, i.e. a particle subject to a force $F(x,t)=\gamma\,{\rm d}_{t}x+\beta+\sqrt{2\alpha}\,\eta(t)$ with negligible inertial term $m\,{\rm d}_{t}^{2}x$. Here, $\eta$ is an stochastic function usually motivated by random collisions.
Thus, equation (11) describes the time evolution of the probability density function of a particle under the influence of random and drag forces, described by $\alpha$ and $\beta$, respectively. In a social mechanical sense, this model describes a population subject to a compelling trend $\beta$ and some noise $\alpha$ that scatters agglomerations of beliefs. For a simple case, the result is a non-uniform distribution that moves in time changing both its mean, as well as its variance. Let us see how.
@@
$\alpha=\frac{\omega_{\alpha}}{2}\sin(\omega_{\alpha}t)\,,\qquad\beta=-\omega_{\beta}\sin(\omega_{\beta}t)\,.$ (14)
@@
Admittedly, this is are rather simplified models of the changes in political stances in the USA, but very descriptive initial ones, nonetheless. This models also show that the framework does not explain the nature of the interactions. Although, it may shed some light on certain processes and, more specifically, on effective phenomena.
As mentioned earlier, the Smoluchowski equation (11) governs the probability density of finding a particle in the position $x$ at time $t$ given that it is subject to an overdamped Brownian motion. However, we wish to include considerations proposed in subsection 3.1, namely, inertial terms of the form $m\,{\rm d}_{t}^{2}x+({\rm d_{x}}m)({\rm d}_{t}x)^{2}$. Then, consider the motion of a particle with position-dependent mass subject to a random force $\eta(x,t)$:
$m\,{\rm d}_{t}^{2}x+m^{\prime}({\rm d}_{t}x)^{2}=\eta\,,$ (16)
where $m^{\prime}={\rm d}_{x}m$. This is the stationary Kardar-Parisi-Zhang equation for $x$ and $t$ playing the role of position $x$, instead. It has an associated Fokker-Planck equation, similar to the Smoluchowski equation, that admits a solution of the form [20]:
$f(t)\sim{\rm e}^{-\frac{m}{2D}\int({\rm d}_{t}x)^{2}{\rm d}t}\,,$ (17)
where $D$ is a diffusion coefficient that captures the effects of the random noise. Notably, $f(t)$ does not depend on $m^{\prime}$. Expression (17) gives the probability density for a given path $x(t)$.